OpenStudy (anonymous):

Which expression represents the difference between 2n and 15? A) 2n - 15 B) 15 - 2n C) Both A and B D) Neither A nor B

OpenStudy (anonymous):

http://mathforum.org/library/drmath/view/69177.html

OpenStudy (anonymous):

i think it is both a nad b? is this right?

OpenStudy (anonymous):

@msumner it's vague what they want. Technically, the answer is as follows...\[\text{A if}\ a\gt\frac{15}2\\\text{B if}\ a < \frac{15}2\\\text{C if}\ a = 0\]

OpenStudy (anonymous):

Oops, I mistook the variable for being a. I intended n.

OpenStudy (anonymous):

so its neather A or B

OpenStudy (anonymous):

It's d because of what they said above :)

OpenStudy (anonymous):

thanks!!! i didnt really get it!

OpenStudy (anonymous):

No, I don't think it's safe to say D. The answer is inconclusive. Ask your instructor how they interpret the difference ;-) Normally, the difference is interpreted as non-negative, so \[ |2n-15|=|15-2n|\]

OpenStudy (anonymous):

i have 2 more questions then i will see if D is the answer!

OpenStudy (anonymous):

A, B, and C are not correct because A and B are indecicive, so that only leaves D.

OpenStudy (anonymous):

ok! thanks

OpenStudy (anonymous):

D is the answer!!!! thanks @pedroSoccer!!!! <3

OpenStudy (anonymous):

No problem :)